OpenStudy (anonymous):

Which equation is in standard form and represents a line with slope –1 through the point (–3, 5)?

OpenStudy (solomonzelman):

you know that y=-x because it's given that the slope is -1 you have to find the y-intercept. y=-x+b plug in your point. (5)=-(-3)+b 5=3+b 2=b Knowing that b (or the y-intercept) is =2 y=-x+2

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (solomonzelman):

Anytime, erinulecki !